# Scalable Unit Harmonization in Medical Informatics Using Bi-directional Transformers and Bayesian-Optimized BM25 and Sentence Embedding Retrieval

### v1

@@ 2.1.1 Primary Dataset
The Optum Clinformatics Data Mart (CDM) constitutes our primary evaluation dataset. This comprehensive, de-identified database integrates administrative health claims data from over 84 million individuals across all 50 U.S. states. The resource contains more than 7.5 billion medical and pharmacy claims, documenting healthcare utilization and associated costs. Data elements include member demographics, detailed medical and pharmacy claims, laboratory results, inpatient confinement records, and provider information.
In this study, we used the 2024 Q1, Q2, and Q3 version of the Optum Clinformatics Data Mart. To support the harmonization process, key fields were extracted from the dataset, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics such as minimum, maximum, mean, and standard deviation. This process resulted in approximately 30,000 unique triads (test, sample, unit) requiring standardization. Of these, 17,244 entries were identified in our internal database as candidates for matching.
@@ 2.2.1 Overall Pipeline Structure
The overall system is structured as an end-to-end processing pipeline (Fig. 1) with distinct stages, each designed to address specific challenges in the harmonization process. The pipeline begins with the Download stage, which involves the extraction of data from the relational database management system (RDBMS), focusing on unique combinations of LOINC codes and reported units, accompanied by pre-computed usage statistics. Following this, the Filtering stage removes entries that fall below a predefined frequency threshold. During this phase, the system also enforces unit name normalization and eliminates inconsistencies to ensure data quality and reduce noise in downstream tasks.
The Update stage enriches the filtered dataset by incorporating additional contextual information. This is achieved through inner joins using LOINC codes to retrieve descriptive labels and standardized mappings for column names, enhancing interpretability and facilitating accurate retrieval. Next, the Predictive Modeling stage implements the core harmonization intelligence (detailed in Section 2.2.2). The pipeline concludes with the Manual Validation stage, which allows domain experts to review and confirm harmonization results through a custom-built user interface. This feedback not only ensures clinical safety and accuracy but also contributes to the iterative improvement of the system.
@@ 2.2.2 Predictive Model Architecture
The predictive model architecture (Fig. 2) constitutes the core intelligence of our system and consists of three primary components. The first component is the Hybrid Retrieval Engine, which combines two complementary approaches. The Lexical Retrieval Module implements BM25 algorithm-based traditional term-matching for precise keyword identification and excels at exact and fuzzy matching of laboratory terminology. Complementing this, the Semantic Retrieval Module utilizes sentence embeddings to capture contextual relationships between terminology and can identify semantically equivalent terms even when they share limited lexical overlap.
The second component is the Bayesian Score Optimizer, which adjusts the weights between lexical and semantic scores to maximize retrieval performance. This component uses a Gaussian Process-based optimization procedure to determine the optimal balance between precision and recall for each query context. Importantly, this optimization is performed offline, prior to inference, ensuring that the optimal weights are precomputed and ready for use during query execution.
The third component is the Transformer-based Reranker, a bidirectional transformer model that acts as the final decision layer, evaluating and reordering candidate matches based on deeper contextual understanding. This component addresses cases where the initial retrieval may have missed subtle semantic distinctions, particularly in complex medical terminology where minor variations can signify clinically significant differences.
@@
The retrieval engine is implemented using an Elasticsearch backend, supporting flexible and scalable querying through JSON-based schema definitions which govern the indexing and searching of laboratory unit metadata. The hybrid retrieval strategy combines lexical scoring—based on BM25 metrics—with semantic similarity, computed via cosine distance between sentence embeddings. The final ranking score for each candidate is calculated as a weighted linear combination of these two components, with weights determined by the Bayesian optimizer.
This modular architecture enables both high accuracy and scalability. The retrieval components provide broad coverage of potential matches, while the reranker delivers precision by applying deeper semantic understanding. Each component can be independently optimized and updated as techniques advance, ensuring the system remains state-of-the-art while maintaining backward compatibility with existing data pipelines.
2.2.3 Sentence Embeddings
We employ the pre-trained transformer-based model bijaygurung/stella-en-400Mv5 [5] with a vector size of 1024 to generate dense vector representations for laboratory test descriptions. This model has been trained using Matryoshka Representation Learning (MRL) [6] to produce hierarchical embeddings at multiple granularities.
We also explored fine-tuning the sentence embedding on laboratory-specific terminology using a synthetic dataset of pairs labeled as either equivalent or dissimilar. However, our experiments demonstrated that the general-purpose model significantly outperformed the fine-tuned models.
2.2.4 BM25 Enhancement
We enhanced our retrieval system by incorporating the BM25 algorithm [7], a state-of-the-art probabilistic relevance framework. BM25 offers a more refined modeling of term frequency, along with document length normalization to mitigate frequency bias. The BM25 score for a document $d$ with respect to a query $q$ is computed as:
@@ 2.2.4 BM25 Enhancement
2.2.5 Search and Ranking
@@
2.2.6 Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization to fine-tune the weights assigned to various components of the search process. This approach allows for the systematic exploration of the search parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. The optimization process primarily focused on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions or units) in the retrieval process. By tuning these parameters, the system can adapt to the specific needs of each search context.
@@ 2.2.6 Bayesian Optimization
Finally, the optimization aimed to maximize the Mean Reciprocal Rank (MRR), a widely used performance metric that evaluates the ranking quality of retrieved results. By focusing on MRR, the system ensures that relevant harmonization candidates appear near the top of the result list, enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks. Through this optimization process, we were able to fine-tune the retrieval system to deliver the most relevant results while minimizing errors and inconsistencies.
2.3 Reranker
The reranker component of the system leverages a bidirectional transformer architecture based on RoBERTa [8] to classify and rank candidate harmonizations. Specifically, we employed the RoBERTa-base-PM-M3-Voc-train-longer model, which was pre-trained on PubMed and PMC articles as well as MIMIC-III clinical notes, using a byte-pair encoding vocabulary learned from PubMed with an additional 50,000 training steps. This domain-adapted model provides enhanced performance on biomedical terminology compared to general-purpose language models. We fine-tuned this biomedical RoBERTa variant on a synthetic dataset consisting of 3,135,557 balanced pairs of laboratory test triads (test, sample, unit), with each pair labeled as either equivalent (1) or non-equivalent (0). The classifier was trained to predict the probability of equivalence between input triads, with this probability score subsequently used as the differentiative ranking metric. This architecture enables the reranker to evaluate each candidate by comparing it directly with the input query, ensuring that the most relevant harmonization suggestions are ranked at the top.
The model also includes token handling mechanisms tailored to the unique terminology of laboratory tests. This is critical, as laboratory names and unit expressions often involve highly specialized vocabulary that can vary across different systems and institutions. By integrating these mechanisms, the reranker can more accurately process domain-specific terms and handle variations in naming conventions.
To further improve the model’s ability to differentiate between similar tests, the reranker incorporates contrastive learning objectives. This helps the system refine its distinction between closely related test descriptions, enabling it to make more nuanced decisions when harmonizing terms that appear almost identical but have subtle differences in meaning or context.
The model was trained on a synthetic dataset, which was created by pairing various test descriptions with different levels of dissimilarity between matching and non-matching pairs. This dataset simulates the real-world challenges of harmonizing laboratory data, with both easy and difficult matching cases. To fine-tune the model, cross-entropy loss was applied, allowing the system to learn to classify test pairs accurately across multiple laboratory categories.
An important feature of the system is its ability to allow for incremental improvement over time. As the system is deployed and interacts with real-world data, manually corrected values from operational use are incorporated into the training dataset. This process of continuous learning enables the model to progressively refine its understanding and handling of complex harmonization patterns, ensuring that it becomes more accurate and effective as it is used in practice.
2.4 Performance Evaluation
@@ 3 Results
We evaluated our harmonization approach using a labeled dataset comprising 17,243 queries from the Optum database. All model parameters were optimized using Bayesian optimization on a stratified random subsample of 2,500 queries.
3.1 Comparative Approach Analysis
@@
We conducted three controlled experiments to evaluate the performance of different retrieval paradigms, each designed to assess the strengths of various retrieval techniques.
@@
3.2 Overall Performance Metrics
@@
3.3 Performance at Different Cutoff Thresholds
@@ 4 Discussion
Our hybrid architecture effectively addresses these limitations by leveraging the complementary strengths of both approaches. The performance metrics across different rank cutoffs demonstrate that the system achieves an optimal balance between precision and recall, with over 83% of queries returning the correct result in the top position.

### v2

# Scalable Unit Harmonization in Medical Informatics via Bayesian-Optimized Retrieval and Transformer-Based Re-ranking

@@ 2.1.1 Primary Dataset
The Optum Clinformatics Data Mart (CDM) [5] constitutes our primary evaluation dataset. This comprehensive, de-identified database integrates administrative health claims data from over 84 million individuals across all 50 U.S. states. The resource contains more than 7.5 billion medical and pharmacy claims, documenting healthcare utilization and associated costs. Data elements include member demographics, detailed medical and pharmacy claims, laboratory results, inpatient confinement records, and provider information.
The large scale and heterogeneity of the dataset provide a rigorous testbed for evaluating the effectiveness and scalability of our unit harmonization system. The vast diversity of lab tests, samples, and reported units reflects real-world variability and complexity in clinical data sources, posing significant challenges for accurate and automated harmonization. Furthermore, the inclusion of data from diverse healthcare settings across all U.S. states ensures that the system is evaluated against a wide range of measurement practices and data quality variations.
Given that Optum CDM is extensively used for pharmacoepidemiology, outcomes research, and healthcare analytics, successful harmonization of units within this dataset directly contributes to improving the reliability and consistency of downstream clinical analyses and decision-making. Therefore, demonstrating our system’s performance on this dataset not only validates its technical robustness but also its practical utility in large-scale medical informatics applications.
@@ 2.2.1 Overall Pipeline Structure
The system is organized as an end-to-end processing pipeline (Fig. 1) comprising three main stages: preprocessing, harmonization, and post-processing.
The preprocessing stage begins by retrieving raw data from the source system, Optum Clinformatics Data Mart (2024 Q1, Q2, and Q3). To support downstream harmonization, key fields are extracted, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics (minimum, maximum, mean, and standard deviation). This process yields approximately 30,000 unique (test, sample, unit) triads requiring standardization, of which 17,244 were found to have candidate matches in our internal database.
Next, the extracted data undergoes cleaning and normalization to ensure consistency in structure and values. A validation step then confirms the integrity of the LOINC codes and associated metadata. Finally, the dataset is enriched by mapping valid LOINC codes to standardized labels and attributes, providing essential context for the harmonization phase.
@@
During harmonization, the system constructs tailored queries that generate ranked proposals for matching equivalents. These queries leverage the hybrid retrieval engine presented in this paper, enabling efficient searches for equivalent data elements. The retrieved results are then processed and structured for evaluation, followed by a reranking step that prioritizes candidates based on relevance scores to enhance overall accuracy.
The pipeline concludes with post-processing, where domain experts review the harmonization outcomes through a custom interface to ensure clinical safety and correctness. Finally, validated results are integrated into the target repository, completing the harmonization process.
As illustrated in Fig. 1, this comprehensive workflow balances automated methods with expert oversight to deliver reliable and high-quality data harmonization.
@@ 2.2.2 Predictive Model Architecture
The core intelligence of the system is embodied in the predictive model architecture (Fig. 2), which consists of a Hybrid Retrieval Engine and a Retrieval Reranker.
The Hybrid Retrieval Engine combines two complementary retrieval methods: the Lexical Retrieval Module, which uses the BM25 algorithm for precise keyword matching and excels at exact and fuzzy matches of laboratory terminology; and the Semantic Retrieval Module, which employs sentence embeddings to capture contextual relationships, enabling detection of semantically equivalent terms even when lexical overlap is limited.
A key design feature within this engine is the Bayesian Score Optimizer, which uses Gaussian Process-based optimization to determine the optimal weighting between lexical and semantic scores. This optimization is performed offline before inference, ensuring that precomputed weights guide the weighted combination of retrieval scores during query execution.
@@
The Retrieval Reranker serves as the final decision layer. This bidirectional transformer model [6, 7, 8] reevaluates and reorders candidate matches based on deeper contextual understanding, addressing subtle semantic distinctions critical in complex medical terminology.
The retrieval engine relies on an Elasticsearch backend, supporting scalable and flexible querying via JSON schema definitions that govern indexing and search of laboratory metadata. Candidate ranking scores are calculated as a weighted linear combination of BM25 lexical scores and semantic similarity scores measured by cosine distance between embeddings, with weights established by the Bayesian optimization design embedded within the Hybrid Retrieval Engine.
This modular architecture achieves both high accuracy and scalability. Broad candidate coverage is provided by the hybrid retrieval engine, while the reranker ensures precision through deeper semantic analysis. Each component can be independently optimized and updated, maintaining backward compatibility and allowing incorporation of future advances.
2.2.3 Baseline Retrieval Using BM25
Our retrieval system incorporates as a base the BM25 algorithm [9], a state-of-the-art probabilistic relevance framework. BM25 offers a refined modeling of term frequency, along with document length normalization to mitigate frequency bias. The BM25 score for a document $d$ with respect to a query $q$ is computed as:
@@ 2.2.3 Baseline Retrieval Using BM25
2.2.4 Contextual Retrieval via Sentence Embeddings
Sentence embeddings [10, 11, 12] serve as the semantic backbone of our retrieval system, enabling context-aware matching between queries and candidate records that go beyond lexical overlap. These embeddings are dense vector representations of text, where semantically similar sentences are mapped to nearby points in the embedding space. In the context of laboratory terminology, this is crucial for recognizing variant expressions of the same clinical concept, such as “blood urea nitrogen” and “BUN,” or for disambiguating polysemous terms based on surrounding context.
2.2.5 Hybrid Lexical–Semantic Retrieval and Ranking
@@
2.2.6 Tuning Lexical and Semantic Weights via Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization [13] to fine-tune the weights assigned to various components of the search process. This approach allows for the systematic exploration of the search parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. The optimization process primarily focused on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions or units) in the retrieval process. By tuning these parameters, the system can adapt to the specific needs of each search context.
@@ 2.2.6 Tuning Lexical and Semantic Weights via Bayesian Optimization
Finally, the optimization aimed to maximize the Mean Reciprocal Rank (MRR) [14], a widely used performance metric that evaluates the ranking quality of retrieved results. By focusing on MRR, the system ensures that relevant harmonization candidates appear near the top of the result list, enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks. Through this optimization process, we were able to fine-tune the retrieval system to deliver the most relevant results while minimizing errors and inconsistencies.
2.3 Transformer-Based Reranking of Retrieved Candidates
The reranker module is responsible for ranking candidate harmonizations presented by previous modules by evaluating their semantic compatibility with the input laboratory test. It operates on test triads, each comprising a test name, a sample type, and a measurement unit. For each candidate triad, the reranker computes a compatibility score based on its similarity to the input, enabling the system to prioritize the most plausible harmonization suggestions. This component plays a critical role in filtering out false matches and refining the output quality of the harmonization pipeline.
To address the complexity of laboratory data, which often includes specialized terminology, institution-specific conventions, and heterogeneous representations, the reranker incorporates tokenization and handling mechanisms specifically adapted to biomedical language. This enables the model to more effectively understand and compare test descriptions that differ in surface form but refer to the same concept.
The reranker is further enhanced by contrastive learning objectives that help the model distinguish between closely related yet non-equivalent test triads. This is especially important for laboratory terms that may differ only slightly in wording or structure but convey distinct clinical meanings. By leveraging a pairwise comparison strategy, the reranker can evaluate subtle differences and improve ranking precision.
Importantly, the model is designed to support incremental learning. As the system is used in real-world settings, manually verified harmonizations are incorporated back into the training set. This continuous learning process allows the reranker to evolve over time and adapt to emerging terminology and use cases, improving its long-term performance and generalization capabilities.
2.4 System Performance Evaluation
@@ 3 Results
In this section, we present the results from the different modules that comprise our retrieval system. We begin with the selection of core components, including BM25 parameter settings and sentence embedding models. This is followed by Bayesian optimization of the weighting parameters used in the combined re-ranking module, training of the re-ranker for optimal ranking performance, and finally, a comparison of the overall system performance under the optimized configuration versus simpler baselines.
3.1 BM25 Parameter Selection
We use the default BM25 configuration provided by Elasticsearch for our retrieval experiments. BM25 is a widely adopted probabilistic ranking function that scores documents based on term frequency, inverse document frequency, and document length normalization. Elasticsearch’s implementation includes two key parameters: $k_{1}$, which controls term frequency saturation and is set to 1.2 by default, and $b$, which governs the degree of document length normalization, with a default value of 0.75. These parameters balance the influence of term frequency and document length on the relevance score. We retained the default settings, as they are generally effective across a range of domains and retrieval tasks, and preliminary experiments did not indicate significant performance gains from additional tuning.
3.2 Sentence Embedding Selection
To select an appropriate sentence encoder for our retrieval task involving laboratory tests, samples, and unit combinations, we evaluated several pre-trained models based on their performance in the Massive Text Embedding Benchmark (MTEB), as shown in Table 1. Our primary selection criteria included model size—favoring models with fewer than one billion parameters for efficiency—and the overall mean task score, which reflects performance across a diverse set of semantic tasks. This balance ensured both strong generalization capabilities and practical resource requirements for integration into our system.
Model # Parameters Embedding Dim. Max Tokens Mean (Task)
SFR-Embedding-Mistral 7B 4096 32768 60.90
stella_en_1.5B_v5 1.5B 8960 131072 56.53
NV-Embed-v2 7B 4096 32768 56.29
stella_en_400M_v5 435M 4096 8192 48.32
bge-small-en-v1.5 33M 512 512 43.76
all-MiniLM-L6-v2 22M 384 256 41.39
Table 1: Comparison of embedding models on MTEB Mean (Task) score.
Among the evaluated models meeting the size restriction, stella_en_400M_v5 — a transformer with approximately 435 million parameters trained using Matryoshka Representation Learning (MRL) [15] — demonstrated strong general performance. While this model produces 4096-dimensional embeddings, the Matryoshka training method distributes semantic information hierarchically across the embedding dimensions, allowing effective selection of lower-dimensional subsets without significant loss of information. Leveraging this property, we used 1024-dimensional representations for our experiments, balancing embedding compactness with semantic richness.
In comparison, lighter models such as bge-small-en-v1.5 and all-MiniLM-L6-v2, with 33 million and 22 million parameters respectively, showed lower Mean (Task) scores, reflecting a trade-off between model size and embedding quality.
Our approach uses these embedding models to generate a candidate set of matches based on semantic similarity. The responsibility of fine-grained selection according to the particularities of our application is delegated to a reranker module, which applies domain-specific criteria to reorder the candidates for the final retrieval.
We also explored fine-tuning using a synthetic dataset of sentence pairs labeled as semantically equivalent or dissimilar, focusing on laboratory-specific terminology. However, our experiments showed that general-purpose pre-trained models consistently outperformed these fine-tuned variants, indicating that large-scale general-domain training still provides strong out-of-the-box performance. Future work may revisit fine-tuning using alternative strategies or additional domain supervision.
3.3 Bayesian Optimization of BM25 and Sentence Embedding Parameters
Figure 3: Bayesian optimization convergence showing the progression of MRR optimization over iterations. The plot displays the objective function values (MRR) across optimization iterations.
The Bayesian optimization process, shown in Figure 3, was performed over approximately 120 different parameter configurations using a validation set of 2,500 distinct laboratory codes. The optimization balanced the lexical contributions of the TEST, SAMPLE, and UNIT fields and a semantic component based on cosine similarity. The cosine similarity was clipped at zero using a max(0, cosine similarity) function before being weighted—assigning zero when dissimilar and up to the full optimized weight when similar. The objective function optimized was the Mean Reciprocal Rank (MRR) across the validation set.
3.4 Reranker Implementation and Training
Figure 4: Validation curves of the training process for the transformer-based reranker. The plot shows the progression of validation loss and validation accuracy over training epochs.
The reranker is implemented using a fine-tuned version of RoBERTa-base-PM-M3-Voc-train-longer model [16], a biomedical variant of RoBERTa [8] that was pre-trained on PubMed abstracts, PMC full-text articles, and MIMIC-III clinical notes. This model employs a byte-pair encoding vocabulary derived from PubMed content and was further pre-trained for 50,000 additional steps to improve domain-specific understanding. Its adaptation to clinical and biomedical language makes it particularly suitable for evaluating laboratory test descriptions.
To fine-tune the model for the harmonization task, we constructed a synthetic dataset composed of 3,135,557 labeled pairs of laboratory test triads. Each pair contains two triads, and the label indicates whether the two triads are semantically compatible or not. The dataset includes pairs with varying levels of dissimilarity to simulate real-world ambiguity. Some pairs differ in all three components—test, sample, and unit—while others differ in only one or two. Pairs labeled as compatible share the same meaning across all components but use different lexical expressions, such as synonyms or variant naming conventions. This stratified construction ensures the model learns to recognize not only exact matches but also semantically equivalent formulations. The reranker finetuning process using as training a synthetic dataset is illustrated in Figure 4.
Training is conducted using binary cross-entropy loss, with the model learning to classify whether a given pair of triads represents a valid equivalence. The output probability serves as a compatibility score and is used directly as the ranking metric. Additionally, contrastive learning objectives are incorporated during training to sharpen the model’s ability to differentiate between closely related but non-equivalent pairs. This helps improve the reranker’s performance in borderline cases where semantic distinctions are subtle.
3.5 Performance evaluation
We evaluated our harmonization approach on a labeled dataset of 17,243 queries from the Optum database, employing the optimal parameters identified through the prior Bayesian optimization. In the following sections, we present the results of this performance evaluation.
3.5.1 Comparative Approach Analysis
@@
We conducted three controlled experiments to evaluate the performance of distinct retrieval paradigms, each aimed at highlighting the strengths of specific retrieval techniques.
@@
3.5.2 Overall Performance Metrics
@@
3.5.3 Performance at Different Cutoff Thresholds
@@ 4 Discussion
Our hybrid architecture effectively addresses these limitations by leveraging the complementary strengths of both approaches. The performance metrics across different rank cutoffs demonstrate that the system achieves an optimal balance between precision and recall, with over $83\%$ of queries returning the correct result in the top position.